OpenStudy (anonymous):

7 + 2(4 – x) = 7 – 2(x – 4).

OpenStudy (anonymous):

Order of Operations

OpenStudy (anonymous):

well the equation will become 7 + (8 - x) = 7 - (2x - 8) because of the Rule caleed PEMDAS then you will have to add 15 - x = 15 + 2x 0 = 3x so it will be x = 0

OpenStudy (anonymous):

thank yuh :)

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

i did it wrong

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

7 + (8 - 2x) = 7 - (2x - 8) 15 - 2x = 15 + 2x 0 = 4x x=0

OpenStudy (anonymous):

orry its just that i am a perfectionist i dont like the equation being wrong even though i got the right answer

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

ok